The temperature of the room the bottles were measured in was 21 degrees, the internal pressure of the bottle was 140.00 atm and the volume 30.0L. The mass of the gas in the bottle was found to be 7.65kg. Which gas is in the bottle? R=0.0821L
1
Expert's answer
2020-11-23T06:52:27-0500

pV = nRT

140.00 atm x 30.0 L = n x 0.0821 L*atm/(mole*K) x (21+273.15) K

n = 174 moles

n = m/Mr

Mr = m/n = 7650 g / 174 moles = 43.9655 g/mol

Mr(CO2) = 12 + 16x2 = 44 g/mol

CO2